a) Calculate the pH of 1.54 MCH3CO2H?

b) Calculate the concentration of H3O+ in 1.54 MCH3CO2H?

c) Calculate the concentration of CH3CO−2 in 1.54 MCH3CO2H?

d) Calculate the concentration of CH3CO2H in 1.54 MCH3CO2H?

e) Calculate the concentration of OH− in 1.54 MCH3CO2H?

f) Calculate the pH of 0.0229 MCH3CO2H?

g) Calculate the concentration of H3O+ in 0.0229 MCH3CO2H?

h) Calculate the concentration of CH3CO−2 in 0.0229 MCH3CO2H?

i) Calculate the concentration of CH3CO2H in 0.0229 MCH3CO2H?

j) Calculate the concentration of OH− in 0.0229 MCH3CO2H?
Answer #1

I have submitted the first four sub-parts, please post multiple question to get the remaining answers

            a) CH_3COOH \rightleftharpoons CH_3COO^- + H^+

Initial              1.54M                       0                   0

Final             (1.54-x)M                   xM                xM

Ka of acetic acid = 1.8 * 10^(-5)

Ka = \frac{[CH_3COO^-][H^+]}{[CH_3COOH]}

1.8 * 10^{-5} = \frac{x^2}{1.54-x}

assuming (1.54-x) is approximately equal to 1.54

x^2 = 1.8 * 1.54 * 10^{-5}

x^2 = 27.72 * 10^{-6}

x = 5.264 * 10^{-3}

pH = 3 - log(5.264) = 2.278

b)

[H^+] = [H_3O^+] = 5.264 * 10^{-3}

c)

[H^+] = [CH_3COO^-] = 5.264 * 10^{-3}

d) [CH_3COOH] = 1.54 - 5.264 * 10^{-3} = 1.534 M